a(T)=-2t(squared)+32T+12
the amount of nitrogen dioxide A was present in the air. T is the number of hours after 6 am. Use this function to find the time at which the nitrogen dioxide level was at its maximum. Graph the parabola.

Answer by stanbon(75887)

a(T)=-2t(squared)+32T+12
the amount of nitrogen dioxide A was present in the air. T is the number of hours after 6 am. Use this function to find the time at which the nitrogen dioxide level was at its maximum. Graph the parabola.
---------------------
a(T) = -2T^2+32T+12
----
Quadratic with a = -2 ; b = 32
---
Max occurs when T = -b/(2a) = -32/(2*-2) = 8 hrs
---
Time at max nitrogen dioxide level 6am + 8 hr = 2pm
=======================================================
Cheers,
Stan H.